# FLAIR-HUB: Large-scale Multimodal Dataset for Land Cover and Crop Mapping Anatol Garioud, Sébastien Giordano, Nicolas David, Nicolas Gonthier Institut national de l’information géographique et forestière (IGN), France

### v1

@@ Abstract
{forest} pic dir tree, where level=0directory,, for tree= s sep=0.1cm, l sep=1cm, font= [FLAIR-HUB dataset [DOMAIN_SENSOR_DATATYPE [ROI [PATCH.tif, file]]] [GLOBAL_ALL_MTD [GLOBAL_MODALITY_*.gpkg, file]] ]
@@ II Related Work
Pixels ×$\times 10^{6}$ Resolution Classes Source Resolution Extent (km2) Source
Vaihingen (vaihingen, ) 82 8 cm 6 visual interpretation 8 cm 1 aerial
EuroSAT (helber2019eurosat, ) 110 50 m 10 EU Urban Atlas (UrbanAtlas, ) 10 m 11 059 Sentinel-2
MultiSenGE (MULTISENGE, ) 534 10 m 14 visual interpretation 10 m 57 433 Sentinel-1&-2
Landcovernet (alemohammad2020landcovernet, ) 589 10 m 7 semi-automatic (MODIS (modis, )) 10 m 58 982 Sentinel-2
MiniFrance (minifrance, ) 1 510 50 m 14 EU Urban Atlas (UrbanAtlas, ) 50 cm 53 000 aerial
DynamicEarthNet (toker2022dynamicearthnet, ) 1 889 3 m 7 visual interpretation 3 m 16 986 Sentinel-1&-2, PlanetFusion
LoopNet (Boonpook_ijgi12010014, ) 3 133 30 m 6 visual interpretation 30 m 2 820k Landsat 8
OpenEarthMap (xia2023openearthmap, ) 4 931 25–50 cm 8 visual interpretation 25–50 cm 799 aerial, UAV,, satellite
Five-Billion-Pixels (tong2023enabling, ) 5 000 4 m 24 visual interpretation 4 m 50 000 Gaofen-2
LoveDA (wang2021loveda, ) 6 000 30 cm 7 visual interpretation 30 cm 536 aerial
Dynamic World (brown2022dynamic, ) 6 348 50m 9 semi-automatic / iteratif 10 m 634k Sentinel-2
DeepGlobe (demir2018deepglobe, ) 6 867 50 cm 7 visual interpretation 50 cm 1 717 Wordlview-2/3, GeoEye-1
BigEarthNet (sumbul2021bigearthnet, ) 8 500 100 m 19 semi-automatic (CLC (cover2000corine, )) 10 m 850 k Sentinel-1&-2
@@
CatLC (garcia2022catlc, ) 25 600 1 m 41 visual interpretation 1 m 25 600 aerial, Sentinel-1&-2, Topo
SeasoNet kossmann2022seasonet 63 353 10 m 33 semi-automatic (CLC (cover2000corine, )) 10 m 748k Sentinel-2
@@ II-B Land cover datasets
Numerous land-cover datasets have been introduced to train semantic segmentation methods, see Table I. Existing datasets usually present a trade-off: they either offer high-resolution annotations but cover a small extent (like Vaihingen (vaihingen, )), or provide large-extent coverage but with low-resolution annotations (such as BigEarthNet (sumbul2021bigearthnet, ), SeasoNet kossmann2022seasonet or SEN12MS (SEN12MS, )), some of them are even the output of fully automatic process zhang2022urbanwatch ; SEN12MS . In contrast, FLAIR garioud_flair_2023a and FLAIR-HUB  offer very high-resolution annotations (20 cm) and covers a large portion of the French territory. FLAIR-HUB is equivalent to SEN12MS (SEN12MS, ) in number of pixels but it provides high-quality very high resolution annotation compared to automatic annotation at 10m ground sampling distance.
@@
MunichCrops russwurm2018multi Sentinel-2 48$\times$48 - 4 284 km2 Munich 1 multi-year 2016-2017 137 k 17 LPIS 42 Go
Crop Type Mapping Ghana m2019semantic Sentinel-1 Sentinel-2 Planets 32$\times$32 - - Ghana - 2016 8 937 4-24 Survey 310 Go
CV4A Kenya kerner2020field Sentinel-2 2016$\times$3035 4 $\sim$2 450 km2 Western Kenya 4 2019 $>$3 000 7 Survey 3.5 Go
ZueriCrops turkoglu2021crop Sentinel-2 24$\times$24 28 k 2400 km2 Zurich - Swiss 1 2019 116 k 5-14-48 LPIS FOAG -
Pastis-R garnot2021panoptic Sentinel-1 Sentinel-2 128$\times$128 2433 $\sim$4 000 km2 France 4 2019 124 k 18 LPIS FR 54 Go
Sen4AgriNet sykas2021sen4agrinet Sentinel-2 366$\times$366 225 k All France Catalonia 1 multi-year 2016-2020 42 M 9-158 LPIS FR-Ca 10 To
@@
AgriSen-COG selea2023agrisen TS-SITS* Sentinel-2 366$\times$366 41 000 - 5 Country 5 multi-temporal 2019-2020 $\sim$7 M 11-* LPIS 28 Go (Label)
@@
Sen4Map sharma2024sen4map Sentinel-2 64$\times$64 335 k Europe 335 k 2018 20 LC $\sim$35 Crops Stat Survey LUCA 1.2 To
@@
AI4SmallFarms persello2023ai4smallfarms Sentinel-2 1 000$\times$1 000 62 Vietnam Cambodia 62 t 439 k - image labelling 1.4 Go
AI4Boundaries d2023ai4boundaries Aerial Ortho(1 m) Sentinel-2 512$\times$512 256$\times$256 7 831 Europe 7 countries 7 831 2019 S2-composite 14.8 M - LPIS 38 Go
Fields of the World kerner2024fields Sentinel-2 multi-date 256$\times$256 70.5 k 166 k km2 4 Continents 24 countries $\sim$80 multi-Year S2 - 1 date 1.63 M - FIBOA EuroCrops -
@@
Dataset Number of Modalities Modality Task Pixels ×$\times 10^{9}$
@@
CatLC (garcia2022catlc, ) 4 LC Segmentation 25.6
@@ III Dataset global description and naming conventions
The second level of the directory structure represents the Regions Of Interest (ROI). A ROI is a set of contiguous patches. In FLAIR-HUB, all the patches are spatially aligned, meaning that regardless the modality and its spatial resolution, they represent the same coverage on the ground : 102.4 m×$\times$102.4 m. This choice has been made to have patches of 512$\times$512 pixels at 0.2 m spatial resolution. The concept of ROIs was introduced to facilitate the annotation process by pooling the labelling cost. FLAIR-HUB dataset is composed of 2 822 ROIs. While patches have fixed sizes on the ground, ROIs have variable sizes. To ensure a perfect nesting of the patches in the ROIs, their width and length are multiples of 512 m. On average, a ROI covers 0.90 km2. Table IV provides information on the distribution of ROI sizes. The ROI names (name of the second level folder) is not unique. The ROI identifier corresponds to internal management information that is not important for the dataset, except for the first two characters, which describe the primary and secondary land use (A=Agricultural, F=Forest, N=Natural, U=Urban).
@@
AERIAL-RLT_PAN 256$\times$256 0.4 m Mono-temporal Panchromatic 1 Relative Domain equalization (DN) UInt8 11.15 Go
AERIAL_RGBI 512$\times$512 0.2 m Mono-temporal Multi-spectral 4 Relative Domain equalization (DN) UInt8 232.76 Go
DEM_ELEV 512$\times$512 0.2 m Mono-temporal Multi-channel 2 Absolute Altitude (m) Float32 365.18 Go
SENTINEL1-ASC_TS 10$\times$10 10.24 m Time Series Multi-channel 2 Absolute $\sigma_{0}$ Backscatter (no unit) Float32 16.68 Go
SENTINEL1-DESC_TS 10$\times$10 10.24 m Time Series Multi-channel 2 Absolute $\sigma_{0}$ Backscatter (no unit) Float32 17.96 Go
SENTINEL2_TS 10$\times$10 10.24 m Time Series Multi-spectral 10 Absolute BOA Reflectance (%) UInt16 41.61 Go
SPOT_RGBI 64$\times$64 1.60 m Mono-temporal Multi-spectral 4 Absolute BOA Reflectance (%) UInt16 6.24 Go
AERIAL_LABEL-COSIA 512$\times$512 0.2 m Mono-temporal Mono-channel 1 Absolute Label UInt8 2.20 Go
ALL_LABEL-LPIS 512$\times$512 0.2 m Mono-temporal Multi-channel 3 Absolute Label UInt8 6.35 Go
SENTINEL2_MSK-SC 10$\times$10 10.24 m Time Series Multi-channel 2 Absolute Probability (%) UInt16 8.39 Go
@@ IV-A Mono-temporal and multi-temporal modalities
• AERIAL_RGBI: The AERIAL_RGBI modality was produced using the ORTHO HR® product; a mosaic of all individual images taken during an aerial survey and mapped onto a cartographic coordinate reference system. Different cameras are used for the aerial image acquisitions. This implies different sensors and consequently, different image characteristics. The final ORTHO HR® product has a spatial resolution of 0.20 m (R, G, B and NIR channels). By design, there are no clouds (and therefore no missing data) in the aerial images. Additionally, some radiometric processing methods are applied to obtain the final product. First, radiometric equalization methods are applied to the individual images. Then, a global radiometric correction is carried out on the merged images covering an entire spatial domain to provide a more satisfying colour balance between channels. We consider this radiometric equalization to be relative: within a domain, the radiometric properties are shared (even if the acquisition dates may differ), but there are shifts in radiometry observable between domains (due to both the date of acquisition and the specific radiometric corrections applied). Therefore, the radiometry of R, G, B and NIR images of the ORTHO HR® product cannot be considered as a physical measurement of channel reflectance. This radiometric information is encoded as an unsigned 8-bit integer. Each AERIAL_RGBI patch is structured with a shape of C×H×W where C = 4 channels ans HxW = 512$\times$512 pixels. To take this modality into account more precisely one can use the AERIAL_MTD_RADIO-STATS and the AERIAL_MTD_DATES metadata. AERIAL_MTD_RADIO-STATS provides the mean and standard deviation per patch for the Red, Green, Blue, and Infrared channels. AERIAL_MTD_DATES gives for each patch the date and time of acquisition, the name of the original individual image and the name of the camera used. The temporal distribution of the aerial images can be seen in Figure 1.
• AERIAL-RLT_PAN: Such as the AERIAL_RGBI modality, the AERIAL-RLT_PAN is an orthoimage produced with individual aerial surveys images. The image mosaic is stated to be from the 1950’s, but in reality, the images intersecting the FLAIR-HUB patches date from 1947 to 1965 (see Figure 1). To keep things simple, we included the date 195X in the domain name of these modalities. 4 domains in the dataset correspond to the same department but with different acquisition dates. These double domains share the same associated AERIAL-RLT_PAN domain. As a result, AERIAL-RLT_PAN is the only modality with 70 domains instead of 74. During these years, aerial photography was not conducted at the department level but rather on a much more local scale, without any specified spatial resolution. Consequently, even though the data has been resampled to a pixel size of 0.4 m in this dataset, the actual spatial resolution can varies significantly from one area to another. Thanks to other unpublished metadata, we estimate that the actual spatial resolution can vary by a factor of 3, ranging from 0.4 m to 1.2 m. The AERIAL-RLT_MTD_DATES metadata provides the acquisition date of each patch and the name of the original historical aerial image. All of the images have been acquired in the Panchromatic interval, being sensitive from blue to red wavelengths. Then, intra-domain radiometric equalization is very challenging due to the nature of old film-based images (vignetting). Locally, there can be significant radiometric differences within a single domain. Additionally, statistical equalization techniques are applied, which result in a significant reduction in global contrast. Therefore, handling the radiometry of this modality is particularly challenging. As with the AERIAL_RGBI modality, we consider the calibration to be relative and not corresponding to a physical measurement. The value is encoded as an 8-bit unsigned integer. The AERIAL-RLT_MTD_RADIO-STATS metadata provides mean and standard deviation information of the panchromatic channel of each patch. Each patch is structured with a shape of H×$\times$×W = 256$\times$256 pixels.
• DEM_ELEV: This modality has two channels : the Digital Surface Model (DSM) and the Digital Terrain Model (DTM). The DSM gives the altitude, in meters (absolute calibration), for each pixel. Thanks to dense matching techniques, the DSM is derived from the same aerial survey that is used to produce the AERIAL_RGBI modality. This gives the DSM the same spatial resolution as the AERIAL_RGBI but more importantly prevents temporal shifts and ground cover changes between the two products making them temporally coherent. However, there still are small geometric differences between the two products, because orthoimages are projected on the DTM. Moreover, dense matching techniques are applied automatically, potentially introducing noise and artifacts. In particular, radiometrically homogeneous areas (e.g., parts of the aerial image with little texture) tend to lead to locally false and varying 3D information. Apart from these artifacts, the vertical accuracy of the DSM can be considered to be twice the spatial resolution (0.4 m). The DTM information comes from the RGE ALTI Digital Terrain Model. This product is a national DTM available at a spatial resolution of 1 m. It is constructed from different sources such as dense matching of aerial images, airborne Lidar or, for mountainous areas, from airborne Synthetic Aperture Radar acquisitions. Depending on the source, the vertical accuracy of the RGE ALTI DTM can vary between 0.3 m and 7 m. DTM provides altitude from ground surface, removing buildings, trees, and other objects. The difference between the DSM and the DTM information is then related to the height of buildings or trees. Each patch is therefore sized C×$\times$×H×$\times$×W = 2$\times$512$\times$512. Each channel represents an altitude and is therefore encoded as a single-precision floating-point (Float32).
• SPOT_RGBI: This modality has been produced using images from satellite SPOT 6-7 that are acquired each year to produce the French SPOT annual mosaic. For each FLAIR-HUB patch, we chose to use images from the annual SPOT 6-7 mosaic of the same year as the AERIAL_RGBI images, ensuring temporal proximity between the two modalities (see Figure 1 for the temporal distribution of both modality). However, there can be a gap of several months between the two observations, leading to differences in object appearances in the images (e.g., agriculture, forest) or actual land cover changes. The acquisition date, time, and the name of the original SPOT 6-7 image are provided in the SPOT_MTD_DATES metadata table. Initially the images are distributed at 1.50 m resolution but were resampled in the FLAIR-HUB dataset to 1.60 m to be a multiple of 0.2 m. The images include four spectral channels: Red, Green, Blue, and Infrared. Such as Sentinel-2 images the radiometric information is calibrated to Level-2A bottom-of-the-atmosphere reflectance (absolute calibration, no unit). Mean and standard deviation of the 4 channels are given in the SPOT_MTD_RADIO-STATS metadata file. The reflectance percentage is encoded as a UInt 16 : 0 = 0% reflectance and to 10 000 = 100% reflectance. The shape of each patch is then C×$\times$×H×$\times$×W = 4$\times$64$\times$64.
@@
• SENTINEL-2_TS & _MSK-SC: Yearly acquisitions from the Copernicus Sentinel-2A and Sentinel-2B satellites are provided for each area. The time-series (_TS) data correspond to Level-2A bottom-of-the-atmosphere reflectance. The dataset includes 10 spectral channels (B02, B03, B04, B05, B06, B07, B08, B8A, B11, B12), excluding the atmospheric bands with a 60 m spatial resolution. To ensure consistency with other modalities and fit the spatial extent of patches to a multiple of 0.2 m, the spatial resolution of Sentinel-2 images was resampled to 10.24 m. While the nominal revisit time at the equator is 5 days, the actual length of the time series varies significantly across different areas due to orbits, ground segment data gaps, or acquisition failures. Consequently, the number of acquisitions in the dataset ranges from 20 to 146 as it can be seen on Figure 2. Level-2A data include geophysical masks for snow and cloud cover (_MSK-SC), which are associated with the time-series data to filter out unfavourable acquisition conditions. To reduce the number of files, each dataset patch is structured with a shape of (T×$\times$×C)$\times$×H×$\times$×W, where T represents the acquisition dates stacked in the first dimension. C = 10 for _TS data and C = 2 for _MSK-SC. The SENTINEL2_MTD_DATES metadata is available for analysing time series. For each patch, it provides a dictionary containing the length of the time series and the acquisition date corresponding to each position in the series. This metadata is complemented by SENTINEL2_MTD_RADIO-STATS, which describes, in dictionary form, the means and standard deviations of the 10 bands per date. Pixels with a probability of being cloudy strictly greater than 0 are excluded from the mean and standard deviation calculation. When the number of samples for statistical computation is zero, the value nan is returned. On the contrary to FLAIR dataset garioud_flair_2023a ; FLAIR-2 , we only consider aligned pixels and do not provide context information from Sentinel-2 time series. That means a reduction of about $93$% of pixels from Sentinel-2 per VHR patches.
• SENTINEL-1ASC & -1DESC_TS: Yearly time series from the C-Band Copernicus Sentinel-1A and Sentinel-1B satellites are provided. Ground Range Detected (GRD) products are used in dual-polarization mode (VV and VH). Both ascending (ASC_TS) and descending orbits (DESC_TS) are included separately, as their incidence angles differ significantly. The Sigma nought ($\sigma_{0}$) backscattering coefficient, which represents the normalized radar backscatter intensity of the surface, is calculated for both polarization channels. It provides essential information on surface properties such as roughness, moisture content, and land cover type. No speckle filtering was applied and data averaging results from the GRD product’s equivalent number of looks, which is approximately 4. Similar to Sentinel-2 time-series data, each Sentinel-1 patch is structured with a shape of (T×$\times$×C)$\times$×H×$\times$×W, where T represents the acquisition dates stacked in the first dimension, and C = 2 corresponding to the two polarization channels (VV and VH). Figure 2 illustrates the number of acquisition available for these modalities. To align with the spatial extent requirements and ensure consistency across modalities, the spatial resolution of Sentinel-1 images was resampled to 10.24 m. Just like the metadata describing the Sentinel-2 series, the files SENTINEL1-ASC_MTD_DATES, SENTINEL1-DESC_MTD_DATES, SENTINEL1-ASC_MTD_RADIO-STATS, and SENTINEL1-DESC_MTD_RADIO-STATS describe the lengths, dates, and statistics of the Sentinel-1 series.
@@
• AERIAL_LABEL-COSIA: The AERIAL_LABEL-COSIA supervision consists in determining the land cover at the pixel-level. It is based on photo-interpretation of the AERIAL_RGBI images and has been manually produced by experts. An initial spatial multi-level image segmentation approach PYRAM was applied, simplifying the labelling at the cluster level. We note that this segmentation was not necessarily final, but was modified interactively when deemed appropriate. It was specified that movable objects (e.g., cars, boats) are not to be annotated as such, but to be classified as the underlying cover. For example, a car on an asphalt road is labelled as an impervious surface FLAIR-1 . We explicitly named this supervision AERIAL_LABEL-COSIA because it is both temporally and geometrically fully consistent with the AERIAL_RGBI images. The AERIAL_LABEL-COSIA patches are then shaped as C×$\times$×H×$\times$×W = 512$\times$512.
The land cover classification consists of 19 classes, ranging from 0 to 18. Table VI provides the list of classes and their respective label number and their frequency in terms of pixel count and percentage (for total, train, valid and test partitions). Please note that the class order has changed compared to previous versions of the dataset (FLAIR-1 ; FLAIR-2 ) to ensure that labels start at 0, classes are thematically organized, and classes corresponding to weak labels appear at the end of the nomenclature. Only the first 15 classes are used in the FLAIR-HUB experiments. Classes 0 to 4 correspond to different types of anthropized areas. These are the categories of interest to be used, for example, to monitor soil land take. Classes 5 to 8 represent natural surfaces without agricultural intensive usage. Classes 9 to 11 relate to agricultural areas, while classes 12 to 17 correspond to forested areas. Class 16 (ligneous) and 17 (mixed) can be considered as weak labels. Due to the polygon-based annotation process , the photo-interpreter may sometimes be unable to distinguish between the deciduous and coniferous classes. In such cases, they use the ligneous label. Similarly, if both deciduous and coniferous are present within the same polygon, the mixed code is used. The geometric segmentation was fine enough to ensure that this weak labels remain very rare ($\approx$0.06of the annotated pixels). Finally, class 19 indicates cases where the photo-interpreter could not provide an annotation (e.g., shadows, uncertainty between certain classes). For a more in-depth analysis of the AERIAL_LABEL-COSIA labels, the AERIAL_MTD_LABEL-COSIA metadata file is available. This file provides, for each patch, the histogram of AERIAL_LABEL-COSIA labels (unit: pixels).
@@
The original LPIS classification includes over 230 crop types. To enhance consistency across years and simplify downstream tasks, these classes were harmonized and organized into a three-level hierarchical taxonomy. While this taxonomy diverges from HCAT schneider2023eurocrops at the second and third levels, it remains broadly similar at the first level. The second and third level classes were determined based on crop availability within the ROIs. Crop types with limited spatial extent (e.g., present in $<10$ ROIs or domains) were merged with similar crops based on phenological traits and growing seasons. Certain LPIS classes—those dominated by ligneous vegetation (e.g., pastured woodlands or oak/chestnut groves), aquatic areas (e.g., salt marshes or reed beds), or artificial surfaces (e.g., greenhouses) were grouped under the background class. In contrast, mixed-crop classes were deliberately left unmerged with homogeneous crop classes. These mixed classes are better interpreted as unknown and are therefore excluded from training loss computation. The final taxonomy contains 23, 31, and 46 classes at levels 1, 2, and 3 respectively. These have been rasterized as separate channels within a single TIFF file, with a spatial resolution of 20 cm—matching that of the AERIAL_RGBI modality

### v2

@@ Abstract
{forest}
@@ II Related Work
Pixels $\times 10^{6}$ Resolution Classes Source Resolution Extent (km2) Source
Vaihingen (vaihingen) 82 8 cm 6 visual interpretation 8 cm 1 aerial
EuroSAT (helber2019eurosat) 110 50 m 10 EU Urban Atlas (UrbanAtlas) 10 m 11 059 Sentinel-2
MultiSenGE (MULTISENGE) 534 10 m 14 visual interpretation 10 m 57 433 Sentinel-1&-2
Landcovernet (alemohammad2020landcovernet) 589 10 m 7 semi-automatic (MODIS (modis)) 10 m 58 982 Sentinel-2
MiniFrance (minifrance) 1 510 50 m 14 EU Urban Atlas (UrbanAtlas) 50 cm 53 000 aerial
DynamicEarthNet (toker2022dynamicearthnet) 1 889 3 m 7 visual interpretation 3 m 16 986 Sentinel-1&-2, PlanetFusion
LoopNet (Boonpook_ijgi12010014) 3 133 30 m 6 visual interpretation 30 m 2 820k Landsat 8
OpenEarthMap (xia2023openearthmap) 4 931 25–50 cm 8 visual interpretation 25–50 cm 799 aerial, UAV,, satellite
Five-Billion-Pixels (tong2023enabling) 5 000 4 m 24 visual interpretation 4 m 50 000 Gaofen-2
LoveDA (wang2021loveda) 6 000 30 cm 7 visual interpretation 30 cm 536 aerial
Dynamic World (brown2022dynamic) 6 348 50m 9 semi-automatic / iteratif 10 m 634k Sentinel-2
DeepGlobe (demir2018deepglobe) 6 867 50 cm 7 visual interpretation 50 cm 1 717 Wordlview-2/3, GeoEye-1
BigEarthNet (sumbul2021bigearthnet) 8 500 100 m 19 semi-automatic (CLC (cover2000corine)) 10 m 850 k Sentinel-1&-2
@@
CatLC (garcia2022catlc) 25 600 1 m 41 visual interpretation 1 m 25 600 aerial, Sentinel-1&-2, Topo
SeasoNet kossmann2022seasonet 63 353 10 m 33 semi-automatic (CLC (cover2000corine)) 10 m 748k Sentinel-2
@@ II-B Land cover datasets
Numerous land-cover datasets have been introduced to train semantic segmentation methods, see Table I. Existing datasets usually present a trade-off: they either offer high-resolution annotations but cover a small extent (like Vaihingen (vaihingen)), or provide large-extent coverage but with low-resolution annotations (such as BigEarthNet (sumbul2021bigearthnet), SeasoNet kossmann2022seasonet or SEN12MS (SEN12MS)), some of them are even the output of fully automatic process zhang2022urbanwatch; SEN12MS. In contrast, FLAIR garioud_flair_2023a and FLAIR-HUB  offer very high-resolution annotations (20 cm) and covers a large portion of the French territory. FLAIR-HUB is equivalent to SEN12MS (SEN12MS) in number of pixels but it provides high-quality very high resolution annotation compared to automatic annotation at 10m ground sampling distance.
@@
MunichCrops russwurm2018multi Sentinel-2 $\times48$ - 4 284 km2 Munich 1 multi-year 2016-2017 137 k 17 LPIS 42 Go
Crop Type Mapping Ghana m2019semantic Sentinel-1 Sentinel-2 Planets $\times32$ - - Ghana - 2016 8 937 4-24 Survey 310 Go
[…]
ZueriCrops turkoglu2021crop Sentinel-2 $\times24$ 28 k 2400 km2 Zurich - Swiss 1 2019 116 k 5-14-48 LPIS FOAG -
Pastis-R garnot2021panoptic Sentinel-1 Sentinel-2 $\times128$ 2433 $\sim$4 000 km2 France 4 2019 124 k 18 LPIS FR 54 Go
Sen4AgriNet sykas2021sen4agrinet Sentinel-2 $\times366$ 225 k All France Catalonia 1 multi-year 2016-2020 42 M 9-158 LPIS FR-Ca 10 To
@@
AgriSen-COG selea2023agrisen TS-SITS* Sentinel-2 $\times366$ 41 000 - 5 Country 5 multi-temporal 2019-2020 $\sim$7 M 11-* LPIS 28 Go (Label)
@@
Sen4Map sharma2024sen4map Sentinel-2 $\times64$ 335 k Europe 335 k 2018 20 LC $\sim$35 Crops Stat Survey LUCA 1.2 To
@@
AI4SmallFarms persello2023ai4smallfarms Sentinel-2 $\times1 000$ 62 Vietnam Cambodia 62 t 439 k - image labelling 1.4 Go
AI4Boundaries d2023ai4boundaries Aerial Ortho(1 m) Sentinel-2 $\times512$ $\times256$ 7 831 Europe 7 countries 7 831 2019 S2-composite 14.8 M - LPIS 38 Go
Fields of the World kerner2024fields Sentinel-2 multi-date $\times256$ 70.5 k 166 k km2 4 Continents 24 countries $\sim$80 multi-Year S2 - 1 date 1.63 M - FIBOA EuroCrops -
@@
Dataset Number of Modalities Modality Task Pixels $\times 10^{9}$
@@
CatLC (garcia2022catlc) 4 LC Segmentation 25.6
@@ III Dataset global description and naming conventions
The second level of the directory structure represents the Regions Of Interest (ROI). A ROI is a set of contiguous patches. In FLAIR-HUB, all the patches are spatially aligned, meaning that regardless the modality and its spatial resolution, they represent the same coverage on the ground : $\times102.4 m$. This choice has been made to have patches of $\times512$ pixels at 0.2 m spatial resolution. The concept of ROIs was introduced to facilitate the annotation process by pooling the labelling cost. FLAIR-HUB dataset is composed of 2 822 ROIs. While patches have fixed sizes on the ground, ROIs have variable sizes. To ensure a perfect nesting of the patches in the ROIs, their width and length are multiples of 512 m. On average, a ROI covers 0.90 km2. Table IV provides information on the distribution of ROI sizes. The ROI names (name of the second level folder) is not unique. The ROI identifier corresponds to internal management information that is not important for the dataset, except for the first two characters, which describe the primary and secondary land use (A=Agricultural, F=Forest, N=Natural, U=Urban).
@@
AERIAL-RLT_PAN $\times256$ 0.4 m Mono-temporal Panchromatic 1 Relative Domain equalization (DN) UInt8 11.15 Go
AERIAL_RGBI $\times512$ 0.2 m Mono-temporal Multi-spectral 4 Relative Domain equalization (DN) UInt8 232.76 Go
DEM_ELEV $\times512$ 0.2 m Mono-temporal Multi-channel 2 Absolute Altitude (m) Float32 365.18 Go
SENTINEL1-ASC_TS $\times10$ 10.24 m Time Series Multi-channel 2 Absolute $\sigma_{0}$ Backscatter (no unit) Float32 16.68 Go
SENTINEL1-DESC_TS $\times10$ 10.24 m Time Series Multi-channel 2 Absolute $\sigma_{0}$ Backscatter (no unit) Float32 17.96 Go
SENTINEL2_TS $\times10$ 10.24 m Time Series Multi-spectral 10 Absolute BOA Reflectance (%) UInt16 41.61 Go
SPOT_RGBI $\times64$ 1.60 m Mono-temporal Multi-spectral 4 Absolute BOA Reflectance (%) UInt16 6.24 Go
AERIAL_LABEL-COSIA $\times512$ 0.2 m Mono-temporal Mono-channel 1 Absolute Label UInt8 2.20 Go
ALL_LABEL-LPIS $\times512$ 0.2 m Mono-temporal Multi-channel 3 Absolute Label UInt8 6.35 Go
SENTINEL2_MSK-SC $\times10$ 10.24 m Time Series Multi-channel 2 Absolute Probability (%) UInt16 8.39 Go
@@ IV-A Mono-temporal and multi-temporal modalities
• AERIAL_RGBI: The AERIAL_RGBI modality was produced using the ORTHO HR® product; a mosaic of all individual images taken during an aerial survey and mapped onto a cartographic coordinate reference system. Different cameras are used for the aerial image acquisitions. This implies different sensors and consequently, different image characteristics. The final ORTHO HR® product has a spatial resolution of 0.20 m (R, G, B and NIR channels). By design, there are no clouds (and therefore no missing data) in the aerial images. Additionally, some radiometric processing methods are applied to obtain the final product. First, radiometric equalization methods are applied to the individual images. Then, a global radiometric correction is carried out on the merged images covering an entire spatial domain to provide a more satisfying colour balance between channels. We consider this radiometric equalization to be relative: within a domain, the radiometric properties are shared (even if the acquisition dates may differ), but there are shifts in radiometry observable between domains (due to both the date of acquisition and the specific radiometric corrections applied). Therefore, the radiometry of R, G, B and NIR images of the ORTHO HR® product cannot be considered as a physical measurement of channel reflectance. This radiometric information is encoded as an unsigned 8-bit integer. Each AERIAL_RGBI patch is structured with a shape of C×H×W where C = 4 channels ans HxW = $\times512$ pixels. To take this modality into account more precisely one can use the AERIAL_MTD_RADIO-STATS and the AERIAL_MTD_DATES metadata. AERIAL_MTD_RADIO-STATS provides the mean and standard deviation per patch for the Red, Green, Blue, and Infrared channels. AERIAL_MTD_DATES gives for each patch the date and time of acquisition, the name of the original individual image and the name of the camera used. The temporal distribution of the aerial images can be seen in Figure 1.
[…]
@@
[…]
@@
[…]
The land cover classification consists of 19 classes, ranging from 0 to 18. Table VI provides the list of classes and their respective label number and their frequency in terms of pixel count and percentage (for total, train, valid and test partitions). Please note that the class order has changed compared to previous versions of the dataset (FLAIR-1; FLAIR-2) to ensure that labels start at 0, classes are thematically organized, and classes corresponding to weak labels appear at the end of the nomenclature. Only the first 15 classes are used in the FLAIR-HUB experiments. Classes 0 to 4 correspond to different types of anthropized areas. These are the categories of interest to be used, for example, to monitor soil land take. Classes 5 to 8 represent natural surfaces without agricultural intensive usage. Classes 9 to 11 relate to agricultural areas, while classes 12 to 17 correspond to forested areas. Class 16 (ligneous) and 17 (mixed) can be considered as weak labels. Due to the polygon-based annotation process , the photo-interpreter may sometimes be unable to distinguish between the deciduous and coniferous classes. In such cases, they use the ligneous label. Similarly, if both deciduous and coniferous are present within the same polygon, the mixed code is used. The geometric segmentation was fine enough to ensure that this weak labels remain very rare ($\approx$0.06% of the annotated pixels). Finally, class 19 indicates cases where the photo-interpreter could not provide an annotation (e.g., shadows, uncertainty between certain classes). For a more in-depth analysis of the AERIAL_LABEL-COSIA labels, the AERIAL_MTD_LABEL-COSIA metadata file is available. This file provides, for each patch, the histogram of AERIAL_LABEL-COSIA labels (unit: pixels).
@@
The original LPIS classification includes over 230 crop types. To enhance consistency across years and simplify downstream tasks, these classes were harmonized and organized into a three-level hierarchical taxonomy. While this taxonomy diverges from HCAT schneider2023eurocrops at the second and third levels, it remains broadly similar at the first level. The second and third level classes were determined based on crop availability within the ROIs. Crop types with limited spatial extent (e.g., present in <10<10 ROIs or domains) were merged with similar crops based on phenological traits and growing seasons. Certain LPIS classes—those dominated by ligneous vegetation (e.g., pastured woodlands or oak/chestnut groves), aquatic areas (e.g., salt marshes or reed beds), or artificial surfaces (e.g., greenhouses) were grouped under the background class. In contrast, mixed-crop classes were deliberately left unmerged with homogeneous crop classes. These mixed classes are better interpreted as unknown and are therefore excluded from training loss computation. The final taxonomy contains 23, 31, and 46 classes at levels 1, 2, and 3 respectively. These have been rasterized as separate channels within a single TIFF file, with a spatial resolution of 20 cm—matching that of the AERIAL_RGBI modality